Mostly clean-up today, then Dan and I left to go home. Gary is still on site, nominally assisting ISS. We heard yesterday that our air shipment has been received in Porto. INEGI will deliver the two boxes to the ops center by the time we return in Jan.
rsw01: added Bulgin caps, but realized we needed Binder caps as well. Stole from motes we intended to use elsewhere.
rsw02: fired up and had 3 issues. Can't see GPS (on an Titan DSM – shouldn't "rs G" still work?). Seems to be plugged in with daughter board to front panel properly. RMY at 20m is spiking – about 1 in 10 samples is bad in the brisk winds we had at the time. They are detected (diagnostic flag is "A", not "0"), so can be removed, but still not what we want. This sensor had spikes when tested in the ops center as well, though they had gone away when I took it out of the shipping box. Likely this is a weak transducer. We could replace it, but eventually won't have any spares. Finally, mote#12 would start up, then die (no blinking LEDs). Replacing with a different mote worked fine. We think that this is the second mote this has happened to.
rsw03: brief stop to read Setra serial number (and get berries from the tree)
rsw04: installed 4-component radiometers and soil sensors at darkhorse, though tower is not yet up. This will require a 15m cable to the soil mote.
tse12: back to finish up from last night. Installed soil sensors. Raised the Ubiquiti to have better line of sight to rne02. Found that EC100 again was in ASCII mode, so had to drop the box to reprogram through the USB port. In general, we need to be more rigorous in our prep. (I think this will be my job in Jan.)
I've created a table of the sites in the EOL Perdigao web page, with some comments.
We left the ops center with 4? towers prepped, primarily the tnw07 60m tower for when INEGI returns.